Jazzieshizzle is a small economy with the nominal GDP of $440, real GDP of $80, and the money supply of $55. Calculate the price level and the velocity of money. 


1
Expert's answer
2021-02-16T05:48:13-0500

Velocity of money "\frac{nominal GDP}{money supply}=\frac{440}{55}=8"

Price level "=" money supply "\times" velocity of money "\div" real GDP "=\frac{55\times 8}{80}=5.5"